# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ 3.2 Convergence analysis and complexity results
If the strong feasibility condition assumed by [9] holds, i.e.,
(41) $\exists\,{\mathbf{y}}\in{{\mathrm{dom}}(g)}\text{ such that }g({\mathbf{y}})\leq-2\rho_{g}D^{2},$
we then have 2 with any (big) $B_{g}>0$ when $g$ is convex and 3 with any (big) $B_{g}>0$ when $g$ is weakly convex. Hence, the conditions in the two cases (i)-(ii) of Proposition 3.10 are satisfied.
Next, we consider a finite-sum weakly convex constraint $g$. For this case, we can obtain an ($\epsilon,\epsilon$)-KKT point of problem Eq. P with a weaker assumption than Eq. 41. Specifically, we require the (uniform) Slater-type CQ assumption that is assumed in [18, 30]. The purpose of making this assumption is to ensure $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all $k\geq 0$ so that Theorem 2.8(iii) can be applied.
Assumption 5 ((uniform) Slater-type CQ).
There exist $B_{g}>~{}0$, $B>0$, and $\overline{\rho}>\rho_{g}\geq 0$, such that for any ${\mathbf{x}}$ satisfying $g({\mathbf{x}})\leq B_{g}$, it holds $g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for some $\mathbf{y}\in{\mathrm{dom}}(g)$.
Remark 3.13.
When $g$ is convex, [18] assumes 1(a) and the existence of ${\mathbf{y}}_{\mathrm{feas}}$ such that $g({\mathbf{y}}_{\mathrm{feas}})<0$. These assumptions naturally lead to 5, with the parameters given by ${\mathbf{y}}={\mathbf{y}}_{\mathrm{feas}}$, $\overline{\rho}=\frac{-g({\mathbf{y}}_{\mathrm{feas}})}{D^{2}}$, $B=\frac{-g({\mathbf{y}}_{\mathrm{feas}})}{2}$, and any (big) $B_{g}>0$. When $g$ is convex, 5 implies 2 with the same value of $B_{g}$, and when $g$ is $\rho_{g}$-weakly convex with $\rho_{g}\geq 0$, it also implies 3 with the same value of $B_{g}$. We will verify these two assertions in Appendix A. Moreover, 5 is clearly implied by the strong feasibility condition Eq. 41.
Before giving our complexity result in this case, we establish a lower bound relating $g_{+}({\mathbf{x}})$ to the distance from ${\mathbf{x}}$ to the feasible set $\Omega$ in the following lemma. Its proof is given in Lemma A.5.
Lemma 3.14.
Suppose that $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, and 1 and 5 hold. If $\textnormal{dist}({\mathbf{x}},\Omega)\leq\min\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\}$, then $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\gamma=\sqrt{B(\overline{\rho}-\rho_{g})/2}$.
In the next lemma, we establish that Algorithm 1 guarantees $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all iterations.
Lemma 3.15.
Suppose that $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, and 14, and 5 hold. In Algorithm 1, let $u^{(k)}=g({\mathbf{x}}^{(k)})$, i.e., ${\mathcal{B}}_{k}=\Xi$ for all $k\geq 0$ in Eq. 25, and choose $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$. Suppose $\|\boldsymbol{\zeta}_{f}^{(k)}\|\leq l_{f}$ for all $k\geq 0$. Set
[…]
(43) $0<\alpha_{k}\leq\min\left\{\frac{C_{0}}{\beta l_{g}^{2}},{\frac{1}{\beta\rho_{g}},}\frac{\gamma}{2l_{f}\rho_{g}},\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}},\,\frac{B_{g}}{4l_{f}l_{g}}\right\},$
where $\gamma=\sqrt{B(\overline{\rho}-\rho_{g})/2}$. If the initial point ${\mathbf{x}}^{(0)}$ satisfies $\textnormal{dist}({\mathbf{x}}^{(0)},\Omega)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$, then it holds $\textnormal{dist}({\mathbf{x}}^{(k)},\Omega)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ and $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all $k\geq 0$.
Proof 3.16.
We prove the desired results by induction on $k$. The claim holds for $k=0$, by the assumption on ${\mathbf{x}}^{(0)}$. Suppose it holds up to iteration $k$. We show that the same bounds hold for $k+1$. We consider the two cases: (i) $g({\mathbf{x}}^{(k)})\geq~{}C_{0}$; (ii) $g({\mathbf{x}}^{(k)})<~{}C_{0}$.
(i) In this case, since $\nu\leq C_{0}\leq g({\mathbf{x}}^{(k)})$, the updating rule from Eq. 1 becomes
${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\boldsymbol{\zeta}_{g}^{(k)}\right)=\left({\mathbf{x}}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}\right)-\alpha_{k}\boldsymbol{\zeta}_{f}^{(k)}.$
Since $0<g({\mathbf{x}}^{(k)})\leq l_{g}\textnormal{dist}\left(\mathbf{x}^{(k)},\Omega\right)<B_{g}$, it follows from Lemma A.1 that $\boldsymbol{\zeta}_{g}^{(k)}\neq\mathbf{0}$. For simplicity of notation, let ${\mathbf{y}}^{(k)}={\mathbf{x}}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}$ and $\mathbf{x}^{\dagger(k)}=\operatorname{Proj}_{\Omega}(\mathbf{x}^{(k)})$. It holds
$\small{\textnormal{dist}^{2}({{\mathbf{y}}}^{(k)},\Omega)\leq\left\|{{\mathbf{y}}}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}=\left\|\mathbf{x}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}}$
$=\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)-2\alpha_{k}\beta\left\langle\boldsymbol{\zeta}_{g}^{(k)},\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\rangle+(\alpha_{k}\beta)^{2}\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}}$
$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+2\alpha_{k}\beta\left(g(\mathbf{x}^{\dagger(k)})-g(\mathbf{x}^{(k)})+\frac{\rho_{g}}{2}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)+(\alpha_{k}\beta)^{2}\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}}$
$=\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-g_{+}(\mathbf{x}^{(k)})+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)+\alpha_{k}\beta\left(\alpha_{k}\beta\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}-g(\mathbf{x}^{(k)})\right)}$
$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-g_{+}(\mathbf{x}^{(k)})+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)}$
$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-{\gamma}{}+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|\right)\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|}$
(44) $\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)-\alpha_{k}\beta\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2},}$
where the third equality comes from $g(\mathbf{x}^{\dagger(k)})=0$ and $g(\mathbf{x}^{(k)})=g_{+}(\mathbf{x}^{(k)})$, the second inequality holds by the $\rho_{g}$-weak convexity of $g$, the third one uses $g(\mathbf{x}^{(k)})\geq C_{0}$, $\|\boldsymbol{\zeta}_{g}^{(k)}\|\leq l_{g}$, and $0<\alpha_{k}\beta\leq\frac{C_{0}}{l_{g}^{2}}$, the fourth one follows by Lemma 3.14, and the last one holds because $-{\gamma}{}\leq-2\rho_{g}\textnormal{dist}({\mathbf{x}}^{(k)},\Omega)=-2\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|$ by induction. Then it follows that
$\vspace{-1mm}\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)$ $\leq\textnormal{dist}({{\mathbf{y}}}^{(k)},\Omega)+\alpha_{k}l_{f}\leq\sqrt{1-\alpha_{k}\beta\rho_{g}}\textnormal{dist}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}l_{f}$
$\leq\textnormal{dist}(\mathbf{x}^{(k)},\Omega),$
where the second inequality uses Eq. 44, and the last one follows from $\beta\geq\frac{2l_{f}l_{g}}{\rho_{g}C_{0}}$ and $\textnormal{dist}\left(\mathbf{x}^{(k)},\Omega\right)\geq\frac{g({\mathbf{x}}^{(k)})}{l_{g}}\geq~{}\frac{C_{0}}{l_{g}}$. Thus $\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ by induction, and $g({{\mathbf{x}}}^{(k+1)})\leq l_{g}\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\frac{B_{g}}{2}$.
(ii) In the second case, we have $g({\mathbf{x}}^{(k)})\leq C_{0}.$ From Eq. 1, we have
$\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)\leq$ $\textnormal{dist}({{\mathbf{x}}}^{(k)},\Omega)+\alpha_{k}(l_{f}+\beta l_{g})$
(45) $\leq$ $\min\left\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\right\}+\alpha_{k}l_{f}+\frac{C_{0}}{l_{g}}\leq\min\left\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\right\},$
where the second inequality follows from $\alpha_{k}\leq\frac{C_{0}}{\beta l_{g}^{2}}$ and $\textnormal{dist}\left({{\mathbf{x}}}^{(k)},\Omega\right)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ by induction hypothesis, and the last one is derived using $\alpha_{k}\leq\min\{\frac{\gamma}{2l_{f}\rho_{g}},\frac{B_{g}}{4l_{f}l_{g}}\}$, and $C_{0}\leq\min\{\frac{\gamma l_{g}}{\rho_{g}},\frac{B_{g}}{4}\}$. We then obtain
$\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)\leq\frac{g({\mathbf{x}}^{(k+1)})}{\gamma}\leq\frac{1}{\gamma}\left(g({\mathbf{x}}^{(k)})+l_{g}\|{{\mathbf{x}}}^{(k+1)}-{{\mathbf{x}}}^{(k)}\|\right)$
(46) $\leq$ $\frac{1}{\gamma}\left(g({\mathbf{x}}^{(k)})+\alpha_{k}l_{g}(l_{f}+\beta l_{g})\right)\leq\frac{C_{0}}{\gamma}+\frac{1}{\gamma}\left(\alpha_{k}l_{g}l_{f}+{C_{0}}{}\right)\leq\min\left\{\frac{\gamma}{2\rho_{g}},\,\frac{B_{g}}{2l_{g}}\right\},$
where the first inequality uses Lemma 3.14 by Eq. 45, the second/third one comes from the updating rule Eq. 1 and 1(c), the fourth one holds because $g({\mathbf{x}}^{(k)})\leq C_{0}$ and $\alpha_{k}\leq\frac{C_{0}}{\beta l_{g}^{2}}$, and the last one follows by $C_{0}\,\leq\,\min\{\frac{\gamma^{2}}{8\rho_{g}},\,\frac{\gamma B_{g}}{8l_{g}}\},$ and $\alpha_{k}\,\leq\,\min\{\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}}\}.$ Hence $g({{\mathbf{x}}}^{(k+1)})\leq l_{g}\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\frac{B_{g}}{2}$.
Finally, combining the two above cases and applying the induction hypothesis, we complete the proof.
Now we are ready to show the complexity result for the finite-sum weakly convex case under 5.
Proposition 3.17.
Suppose $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, and 14, and 5 hold. In Algorithm 1, let $u^{(k)}=g({\mathbf{x}}^{(k)})$, i.e., ${\mathcal{B}}_{k}=\Xi$ for all $k\geq 0$ in Eq. 25, and choose $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$. Suppose $\|\boldsymbol{\zeta}_{f}^{(k)}\|\leq l_{f}$ for all $k\geq 0$. Then given $\epsilon>0$Algorithm 1 can produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation with $O(\epsilon^{-4})$ stochastic subgradients of $f$, and $O(\epsilon^{-4})$ deterministic subgradients and function evaluations of $g$.
Proof 3.18.
Choose $\overline{\epsilon}$ and $\beta$ to satisfy the conditions in Eq. 17 and Eq. 42 and $\beta>\frac{2B_{f}}{B_{g}}$, and let $\nu$ satisfy $0<\nu\leq\min\{C_{0},{B}_{g},\frac{\epsilon}{\beta},\frac{\epsilon}{2}\},$ where $C_{0}$ is given in Lemma 3.15. Also, choose the stepsize as
$\alpha_{k}=\alpha=\min\left\{\frac{C_{0}}{\beta l_{g}^{2}},{\frac{1}{\beta\rho_{g}},}\frac{\gamma}{2l_{f}\rho_{g}},\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}},\,\frac{B_{g}}{4l_{f}l_{g}},\,\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta l_{g}L_{g}L_{F}}\right\}.\vspace{-2mm}$
Since $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$ and $u^{(k)}=g({\mathbf{x}}^{(k)})$, it holds $\sigma=\sigma_{2}=0$. Hence, we can still obtain Eq. 38 from Eq. 40 by setting $q=1$, $S_{2}=N$, (i.e., taking the whole data), and $K=\lceil 16\overline{\epsilon}^{-2}\alpha^{-1}\Delta\rceil$, thus a near-$\overline{\epsilon}$ stationary point ${\mathbf{x}}^{*}$ of Eq. PP in expectation can be produced within $K$ iterations. Now by Theorem 2.8(iii) and Lemma 3.15, ${\mathbf{x}}^{*}$ is also an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation. It takes $K=O(\epsilon^{-4})$ stochastic subgradient of $f$ and $K=O(\epsilon^{-4})$ deterministic subgradients and function evaluations of $g$. The proof is then completed.
@@ 4 Numerical experiments
In this section, we evaluate the empirical performance of the proposed Algorithm 1 on two fairness-constrained problems, which are instances of Eq. P. We compare its performance with three state-of-the-art approaches, including two different double-loop IPP [9, 30] and the SSG in [18]. All the experiments are performed on iMac with Apple M1, a 3.2 GHz 8-core processor, and 16GB of RAM running MATLAB R2024a.
4.1 Datasets and tested problems
The experiments are conducted on two datasets: a9a [20], and COMPAS [5]. The a9a dataset contains 48,842 data points with 123 features, and is used to predict income levels; gender is the protected attribute, and fairness is assessed between males and females. The COMPAS dataset includes 6,172 data points and 16 features, focused on predicting recidivism risk, with fairness evaluated between Caucasian and non-Caucasian groups. We split each dataset into two subsets with a ratio of 2:1. The larger set is used as $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, and the smaller set is further split into a protected set $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and an unprotected set $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$.
Problem 1. The first tested problem is a classification problem with ROC-based fairness [37]. It is formulated in Eq. 4. For this problem, we have $\rho_{g}=0$ and $\rho_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}$. That is, $\Phi(\cdot)-\Phi^{*}-\kappa_{1}$ is convex. We obtain the minimum $\Phi^{*}$ and a solution ${\mathbf{x}}^{*}$ by applying a stochastic subgradient method to a large number of iterations on $\min_{{\mathbf{x}}}\Phi({\mathbf{x}})$. Afterward, we set $\kappa_{1}=0.001\cdot\Phi^{*}$, $D=5\|{\mathbf{x}}^{*}\|$, and let $\Theta$ consist of 400 points equally spaced between $\uline{{\mathbf{a}}}-0.5\left(\overline{{\mathbf{a}}}-\uline{{\mathbf{a}}}\right)$ and $\overline{{\mathbf{a}}}+0.5\left(\overline{{\mathbf{a}}}-\uline{{\mathbf{a}}}\right)$, where $\overline{{\mathbf{a}}}=\max_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$ and $\uline{{\mathbf{a}}}=\min_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$. For all the compared methods, we initialize ${\mathbf{x}}^{(0)}={\mathbf{x}}^{*}$.

### v2

@@ 3.3 Complexity result
When $\mathbf{g}$} is deterministic, our algorithm still converges to an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation without requiring the condition $B_{g}>m\rho_{g}D^{2}$. This condition appears only once in our analysis, specifically in the proof of inequality (22), to ensure well-definedness of parameters $\beta$ and $\nu$ in (17). Under this assumption, inequality (22) and the subsequent arguments guarantee that $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$, which then enables us to use the Slater-type CQ assumption. In contrast, for the deterministic case, the constraints $g_{i}({\mathbf{x}})\leq 0,\forall\,i\in[m]$ can be equivalently expressed as $\max_{i\in[m]}g_{i}({\mathbf{x}})\leq 0$. Leveraging this reformulation, we directly utilize the established convergence result for the special case of $m=1$, detailed in our arXiv first version [44]. This prior result ensures that, provided the initial point is sufficiently close to the feasible set $\mathcal{X}$}, all iterates produced by our method satisfy $\max_{i\in[m]}g_{i}(\widehat{{\mathbf{x}}}^{(k)})\leq B_{g}$. Consequently, the assumption $B_{g}>m\rho_{g}D^{2}$ becomes unnecessary for this case, provided that we can obtain a (near) feasible initial point.
@@ 4 Numerical experiments
In this section, we evaluate the empirical performance of the proposed Algorithm 1 and compare to three state-of-the-art approaches, including two different IPPs [10, 46] and the SSG in [27]. All the experiments are performed on iMac with Apple M1, a 3.2 GHz 8-core processor, and 16GB of RAM running MATLAB R2024a.
4.1 Tested problems and Datasets
We conduct experiments on solving two fairness-constrained problems and one Neyman-Pearson classification problem, which are instances of Eq. P.
Problem 1. The first tested problem is a classification problem with ROC-based fairness [56]. It is formulated in Eq. 5. We obtain the minimum $\Phi^{*}$ and a solution $\mathbf{x}$$^{*}$ by applying a stochastic subgradient method to a large number of iterations on $\min_{{\mathbf{x}}\in{\mathcal{X}}}\Phi({\mathbf{x}})$. Afterward, we set $\kappa_{1}=0.001\cdot\Phi^{*}$, $D=5\|{\mathbf{x}}^{*}\|$, and let $\Theta$ consist of 400 points equally spaced between $\uline{a}-0.5\left(\overline{a}-\uline{a}\right)$ and $\overline{a}+0.5\left(\overline{a}-\uline{a}\right)$, where $\overline{a}=\max_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$ and $\uline{a}=\min_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$. For all the compared methods, we initialize ${\mathbf{x}}^{(0)}={\mathbf{x}}^{*}$. For this problem, we have $l_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|$, $l_{g}=\frac{1}{n}\sum_{i=1}^{n}\|{\mathbf{a}}_{i}\|$, $\rho_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}$, and $\rho_{g}=0$ from [27]. That is, $\Phi(\cdot)-\Phi^{*}-\kappa_{1}$ is convex, and 1 is satisfied. Moreover, 2 holds with parameters $B_{g}=1$, $\overline{\rho}=\frac{-\max_{i\in[m]}g_{i}({\mathbf{x}}^{*})}{D^{2}}$, and $B=\frac{-\max_{i\in[m]}g_{i}({\mathbf{x}}^{*})}{2}$.